OpenStudy (anonymous):

divide and simplify (7x-28)/33 / (x-4)/22x

OpenStudy (saifoo.khan):

\[\frac{7(x-4)}{33} \times \frac{22x}{x-4}\]

OpenStudy (anonymous):

i got as far as 154x^2-616x / 3x-12 but not sure if it is right or where to go from there

OpenStudy (saifoo.khan):

\[\frac{7\cancel{(x-4)}}{33} \times \frac{22x}{\cancel{x-4}}\]

OpenStudy (anonymous):

ahhhh, ok saifoo, thank you

OpenStudy (saifoo.khan):

No problem, now cancel out 22/33 = 2/3 So finally you are left with: 7*2x ---- 3

OpenStudy (saifoo.khan):

14x ---- 3 Answer